what is 36 nickels to 37 dimes

Found 2 solutions by ChrisDS, fcabanski:
Answer by ChrisDS(7)

36*5=180 cents
180/10= 18 dimes

Answer by fcabanski(1391)

This is a ratio 36N / 37D. A nickle is .05 and a dime is .1


36*.05/37*.1 = 1.80/3.70 which simplifies to 1.8/3.7 or 1.8 : 3.7